The sum of the third and tenth term of an AP is 12. Given that the sum of the 6th terms is 3 less than one third of the eight term. Find the first term and the common difference.


1
Expert's answer
2021-09-23T16:43:04-0400

Let a be first term and d- difference of arithmetic  progression

Then "a_3=a+2\cdot d" - the third term

"a_{10}=a+9\cdot d" - the tenth term

Therefore from the first condition we have an eqution^

"a+2d+a+9d=2a+11d=12"

The sum of 6th terms is "a_1+a_2+a_3+a_4+a_5+a_6=\frac{a_1+a_6}{2}\cdot 6="

"=(a_1+a_6)\cdot 3=(2a+5d)\cdot 3=6a+15d"

So second condition can be written as "6a+15d=\frac{a_8}{3}-3=\frac{a+7d}{3}-3"

or as "17a+38d= -9"

Now we have the system:

"\begin{cases}
 2a+11d=12\\
 17a+38d= -9
\end{cases}"

From the first eqution we have "a=6-\frac{11}{2}\cdot d"

After we insert a in second eqution

"17a+38d=17\left( 6-\frac{11}{2}\cdot d \right)+38\cdot d=-9\\
102-\frac{187}{2}\cdot d+38\cdot d=-9;\\
204-187\cdot d+76\cdot d=-18;\\
222=111\cdot d;
d=\frac{222}{111}=2;
a=6-\frac{11}{2}\cdot 2=6-11=-5;\\"

So answer is:

a=-5- the first term of AP;

d=2- difference of AP;